# DrivingForward: Feed-forward 3D Gaussian Splatting for Driving Scene Reconstruction from Flexible Surround-view Input

### v1

@@ 1 Introduction
3D scene reconstruction is critical for understanding driving scenes. Modern self-driving assistance cars are usually equipped with several cameras to capture surrounding scenes. Real-time reconstruction of driving scenes from sparse vehicle-mounted cameras contributes to various downstream tasks in autonomous driving, including online mapping (Li et al. 2022a), BEV perception (Li et al. 2022b; Liang et al. 2022; Liu et al. 2023), and 3D detection (Chen et al. 2023; Cai et al. 2023). However, the real-time computing required by downstream tasks and the sparse surrounding views challenge the driving scene reconstruction.
Neural Radiance Fields (NeRF) (Mildenhall et al. 2021) and 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) have significantly progressed the development of 3D scene reconstruction. DrivingGaussian (Zhou et al. 2024), StreetGaussian (Yan et al. 2024), and AutoSplat (Khan et al. 2024) further explore the reconstruction of driving scenes. While these methods demonstrate strong capability in novel view synthesis, they are scene-optimized methods that require dozens of images and expensive computing time to reconstruct just one scene. These offline reconstruction methods are unsuitable for real-time downstream tasks in autonomous driving, thereby limiting their practicality.
@@
  • To our knowledge, DrivingForward is the first feed-forward Gaussian Splatting model for driving scenes. It achieves real-time reconstruction from sparse vehicle-mounted cameras and supports flexible multi-frame inputs of surrounding views .
@@ 2.2 Driving Scene Reconstruction
Based on NeRFs or 3DGS, a few methods (Guo et al. 2023; Yang et al. 2023; Zhou et al. 2024; Yan et al. 2024; Khan et al. 2024) extend reconstruction specifically for driving scenes. However, most of them focus on accurate 3D or 4D reconstruction for a single scene. These offline reconstruction methods limit the application of real-time downstream tasks in driving scenes. In contrast, our method is designed for real-time reconstruction to support online downstream tasks. A few NeRFs-related methods (Huang et al. 2024; Yang et al. 2024) also learn 3D representations for online driving, while they mainly focus on other downstream tasks instead of reconstruction. Another related work DistillNeRF (Wang et al. 2024) proposes a generalizable model for driving scenes and achieves reasonable novel-view synthesis. However, it requires a pre-training stage that involves offline optimization of NeRFs for each scene, which is extremely costly and necessitates 8 A100 GPUs (8 × 80GB). In contrast, our DrivingForward only requires a single GPU with 48GB. DistillNeRF also relies on LiDAR for training, while our method is trained from RGB images alone. Despite DitillNeRF requiring more training resources and additional LiDAR data, it achieves only suboptimal performance compared to our DrivingForward under the same setting.
@@ 3.2 Scale-aware Localization
$\mathcal{P}(I_{i}^{t},I_{i}^{t^{\prime}})$ $\rightarrow T_{i}^{t\rightarrow t^{\prime}}$ (1)
@@
$I_{src}=\left\{\begin{array}[]{ll}I_{i}^{t^{\prime}},t^{\prime}\in\{t+1,t-1\}&\text{for temporal}\\ I_{j}^{t},&\text{for spatial}\\ I_{j}^{t^{\prime}},t^{\prime}\in\{t+1,t-1\}&\text{for spatial-temporal}\end{array}\right.$ (5)
@@
$\begin{array}[]{c}T^{trg\rightarrow src}=\\ \left\{\begin{array}[]{ll}T_{i}^{t\rightarrow t^{\prime}},t^{\prime}\in\{t+1,t-1\}&\text{for temporal}\\ E_{j}E_{i}^{-1}&\text{for spatial}\\ E_{j}E_{i}^{-1}T_{i}^{t\rightarrow t^{\prime}},t^{\prime}\in\{t+1,t-1\}&\text{for spatial-temp.},\\ \end{array}\right.\end{array}$ (6)
@@ 3.4 Joint Training Strategy
$\mathcal{L}_{render}=\beta L_{2}+\gamma L_{LPIPS}$ (12)
@@
EmerNeRF Arxiv 2023 SF 114 $\times$ 228 20.95 0.585 -
DistillNeRF Arxiv 2024 SF 114 $\times$ 228 20.78 0.590 -
Ours - SF 352 $\times$ 640 21.67 0.727 0.259
Ours - SF 114 $\times$ 228 21.76 0.767 0.194
@@ 4.1 Experimental Setup
For feed-forward methods, we compare our model with MVSplat (Chen et al. 2024), pixelSplat (Charatan et al. 2024), and DistllNeRF (Wang et al. 2024) (along with its comparison methods (Yang et al. 2023, 2024; Huang et al. 2024)). MVSplat and pixelSplat are designed for training on datasets with densely overlapping inputs. Given that temporally adjacent frames have significantly more overlap than spatially adjacent frames, we use MF mode to meet their overlapping requirements. We retrained them by combining their publicly available codebases with our dataset and data loader. Further training details are present in the supplementary material. DistillNeRF does not release the code and cannot be trained. To enable a fair comparison, we align our method with the settings specified in its paper, including input, rendered novel view, image resolution, and validation scenes. Since SF mode aligns with DistillNeRF’s input and novel views, we train our model using this mode.
@@ 4.5 Ablation Studies
To validate the effectiveness of the joint training strategy, we first train the scale-aware localization module, which includes only the depth and pose networks. We then use this pre-trained model to train the Gaussian network (2nd row, w/o Joint Training). The result in Table 4 shows that joint training enables end-to-end training without decreasing performance and even gains a better performance. The improvement indicates that the supervision of rendering images also contributes to the previous localization module.
In addition to the estimated depth map, the shared image feature from the depth network is another key component for connecting to the subsequent Gaussian network. To demonstrate the importance of this connection, we ablate the shared image feature and instead construct another feature extractor like the depth network encoder for the Gaussian network. As illustrated in Table 4, the model “w/o Image Feature Share” (3rd row) exhibits a drop of more than 2 dB PSNR. This highlights the importance of the shared image feature in connecting the positions and other parameters of Gaussian primitives.
@@ 5 Conclusion
We introduce DrivingForward, a feed-forward Gaussian Splatting model that achieves real-time driving scene reconstruction from flexible surround-view input. To solve the problem of minimal overlap of the surrounding views, we predict the Gaussian primitives from each image independently and propose the scale-aware localization to obtain the multi-view consistent position for Gaussian primitives. A Gaussian network predicts primitives’ other parameters from each image. The individual prediction enables flexible multi-frame inputs of surrounding views. Our method does not require depth ground truth and is extrinsic-free during training. At the inference stage, our model is faster than other methods and achieves higher reconstruction quality for driving scenes compared with both existing feed-forward and scene-optimized reconstruction methods.

### v2

@@ 1 Introduction
3D scene reconstruction is critical for understanding driving scenes. Modern self-driving assistance cars are usually equipped with several cameras to capture surrounding scenes. Real-time reconstruction of driving scenes from sparse vehicle-mounted cameras contributes to various downstream tasks in autonomous driving, including online mapping (Li et al. 2022a), BEV perception (Li et al. 2022b; Liang et al. 2022; Liu et al. 2023), scene understanding (Ji et al. 2024; Sun et al. 2024a; Tan et al. 2023; Sun et al. 2024b; Ma et al. 2024) and 3D detection (Chen et al. 2023; Cai et al. 2023). However, the real-time computing required by downstream tasks and the sparse surrounding views challenge the driving scene reconstruction.
Neural Radiance Fields (NeRF) (Mildenhall et al. 2021) and 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) have significantly progressed the development of 3D scene reconstruction. DrivingGaussian (Zhou et al. 2024), StreetGaussian (Yan et al. 2024), S3Gaussian (Huang et al. 2024a), and AutoSplat (Khan et al. 2024) further explore the reconstruction of driving scenes. While these methods demonstrate strong capability in novel view synthesis, they are scene-optimized methods that require dozens of images and expensive computing time to reconstruct just one scene. These offline reconstruction methods are unsuitable for real-time downstream tasks in autonomous driving, thereby limiting their practicality.
@@
  • To our knowledge, DrivingForward is the first feed-forward Gaussian Splatting model for surround-view driving scenes. It achieves real-time reconstruction from sparse vehicle-mounted cameras and supports flexible multi-frame inputs of surrounding views.
@@ 2.2 Driving Scene Reconstruction
Based on NeRFs or 3DGS, a few methods (Guo et al. 2023; Yang et al. 2024b; Zhou et al. 2024; Yan et al. 2024; Khan et al. 2024) extend reconstruction specifically for driving scenes. However, most of them focus on accurate 3D or 4D reconstruction for a single scene. These offline reconstruction methods limit the application of real-time downstream tasks in driving scenes. In contrast, our method is designed for real-time reconstruction to support online downstream tasks. A few NeRFs-related methods (Huang et al. 2024b; Yang et al. 2024a) also learn 3D representations for online driving, while they mainly focus on other downstream tasks instead of reconstruction. Another related work DistillNeRF (Wang et al. 2024) proposes a generalizable model for driving scenes and achieves reasonable novel-view synthesis. However, it requires a pre-training stage that involves offline optimization of NeRFs for each scene, which is extremely costly and necessitates 8 A100 GPUs (8 × 80GB). In contrast, our DrivingForward only requires a single GPU with 48GB. DistillNeRF also relies on LiDAR for training, while our method is trained from RGB images alone. Despite DitillNeRF requiring more training resources and additional LiDAR data, it achieves only suboptimal performance compared to our DrivingForward under the same setting.
@@ 3.2 Scale-aware Localization
$\mathcal{P}(I_{i}^{t},I_{i}^{t^{\prime}})$ $\rightarrow T_{i}^{t\rightarrow t^{\prime}},$ (1)
@@
$I_{src}=\left\{\begin{array}[]{ll}I_{i}^{t^{\prime}},t^{\prime}\in\{t+1,t-1\}&\text{for temporal,}\\ I_{j}^{t},&\text{for spatial,}\\ I_{j}^{t^{\prime}},t^{\prime}\in\{t+1,t-1\}&\text{for spatial-temporal,}\end{array}\right.$ (5)
@@
$\begin{array}[]{c}T^{trg\rightarrow src}=\\ \left\{\begin{array}[]{ll}T_{i}^{t\rightarrow t^{\prime}},t^{\prime}\in\{t+1,t-1\}&\text{for temporal,}\\ E_{j}E_{i}^{-1}&\text{for spatial,}\\ E_{j}E_{i}^{-1}T_{i}^{t\rightarrow t^{\prime}},t^{\prime}\in\{t+1,t-1\}&\text{for spatial-temp,}\\ \end{array}\right.\end{array}$ (6)
@@ 3.4 Joint Training Strategy
$\mathcal{L}_{render}=\beta L_{2}+\gamma L_{LPIPS},$ (12)
@@
EmerNeRF ICLR 2024 SF 114 $\times$ 228 20.95 0.585 -
DistillNeRF NeurIPS 2024 SF 114 $\times$ 228 20.78 0.590 -
Ours AAAI 2025 SF 352 $\times$ 640 21.67 0.727 0.259
Ours AAAI 2025 SF 114 $\times$ 228 21.76 0.767 0.194
@@ 4.1 Experimental Setup
For feed-forward methods, we compare our model with MVSplat (Chen et al. 2024), pixelSplat (Charatan et al. 2024), and DistllNeRF (Wang et al. 2024) (along with its comparison methods EmerNeRF (Yang et al. 2024b), Unipad (Yang et al. 2024a), and SelfOcc (Huang et al. 2024b)). MVSplat and pixelSplat are designed for training on datasets with densely overlapping inputs. Given that temporally adjacent frames have significantly more overlap than spatially adjacent frames, we use MF mode to meet their overlapping requirements. We retrained them by combining their publicly available codebases with our dataset and data loader. DistillNeRF does not release the code and cannot be trained. To enable a fair comparison, we align our method with the settings specified in its paper, including input, rendered novel view, image resolution, and validation scenes. Since SF mode aligns with DistillNeRF’s input and novel views, we train our model using this mode.
@@ 4.5 Ablation Studies
To validate the effectiveness of the joint training strategy, we first train the scale-aware localization module, which includes only the depth and pose networks. We then use this pre-trained model to train the Gaussian network (2nd row, w/o Joint Training). The result in Table 4 shows that joint training enables end-to-end training without decreasing performance and even gains a better performance. This indicates that the joint training is able to facilitate model convergence and can lead to improvement.
In addition to the estimated depth map, the shared image feature from the depth network is another key component for connecting to the subsequent Gaussian network. To demonstrate the importance of this connection, we ablate the shared image feature and instead construct another feature extractor like the depth network encoder for the Gaussian network. As illustrated in Table 4, the model “w/o Image Feature Share” (3rd row) drops more than 2 dB PSNR. This highlights the importance of the shared image feature in connecting the positions and other parameters of Gaussian primitives.
@@ 5 Conclusion
We introduce DrivingForward, a feed-forward Gaussian Splatting model that achieves real-time driving scene reconstruction from flexible surround-view input. To solve the problem of minimal overlap of the surrounding views, we predict the Gaussian primitives from each image independently and propose the scale-aware localization to obtain the multi-view consistent position for Gaussian primitives. A Gaussian network predicts primitives’ other parameters from each image. The individual prediction enables flexible multi-frame inputs of surrounding views. Our method does not require depth ground truth and is extrinsic-free during training. At the inference stage, our model is faster than other methods and achieves higher reconstruction quality for driving scenes compared with both existing feed-forward and scene-optimized reconstruction methods. In the future, this work is expected to integrate with human perception research (Guo et al. 2024) to develop a more intelligent human-in-scene perception system.
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China (62302167, U23A20343, 62222602, 62176092, 62472282, 72192821); in part by Shanghai Sailing Program (23YF1410500); in part by Chenguang Program of Shanghai Education Development Foundation and Shanghai Municipal Education Commission (23CGA34); in part by the Fundamental Research Funds for the Central Universities (YG2023QNA35); in part by YuCaiKe [2023] (231111310300).